aglue company needs to make some glue worth $120. it wants to use 150 barrels of glue worth $100 per barrel, along with some glue worth $150 per barrel, and some glue worth $190 per barrel. it must use the same number of barrels worth $150 per barrel and $190 per barrel. how much of the $150 and $190 glue will be needed? how many barrels of $120 glue will be producded?
so fr i have:
x=# of barrels of $100 glue
y=# of barrels of $150 glue
z=# of barrels of $190 glue
x=150
y=z

Answer by stanbon(48568)

aglue company needs to make some glue worth $120. it wants to use 150 barrels of glue worth $100 per barrel, along with some glue worth $150 per barrel, and some glue worth $190 per barrel. it must use the same number of barrels worth $150 per barrel and $190 per barrel. how much of the $150 and $190 glue will be needed? how many barrels of $120 glue will be producded?
so fr i have:
x=# of barrels of $100 glue
y=# of barrels of $150 glue
z=# of barrels of $190 glue
x=150
y=z
----------------------------------------
Value Equation: 100*150 + 150y + 190z = Value
Since y = z, Value = 100*150 + (150+190)y = Value
Value = 120(150+2y)
--------------------------------
Therefore
100*150 + (150+190)y = 120(150+2y)
15000 + 340y = 18000+240y
100y = 3000
y = 30 (amt. of $150 glue in the mix)
z = 30 (amt. of $190 glue in the mix)
x = 150 (amt of $100 glue in the mix)
---
To number of barrels of glue 150 + 60 = 210

============================================
Cheers,
Stan H.